# Robust and Resilient Networks with Integrated Sensing, Communication and Computation

### v1

@@ IV-B Open Problems
Efficient and scalable resource management for ISCC: Although widely used in conventional networks, diversity and redundancy schemes are typically designed for individual purposes, without jointly addressing sensing, computing, and communication. In the context of ISCC networks, new strategies that can efficiently integrate diversity and redundancy across all three components are needed, particularly for tasks that exhibit intrinsic dependencies among these functions. Moreover, the joint optimization of sensing, computing, and communication resources results in inherently high-dimensional and complex optimization problems. Hence, the development of efficient, scalable solution methods is critical to enable practical deployment and operation of robust and resilient ISCC systems.
@@
Task-oriented and criticality-aware ISCC: ISCC networks must handle heterogeneous tasks with varying requirements and urgency levels simultaneously. To enhance robustness through proactive resource management, it is essential that the system is aware of these task-level differences when allocating resources. This awareness becomes even more critical during disruptions, when resource availability is constrained and careful prioritization is essential. One promising approach to address this issue is network slicing, which enables application-aware resource allocation. However, conventional network slicing primarily focuses on communication resources, which is insufficient in the ISCC context. Therefore, holistic slicing approaches that can simultaneously manage sensing, communication, and computation resources are required.
@@
Joint source–channel–computing network coding schemes: To support network-level resilience, coding must extend beyond individual node types and retain the network functionality even when some sensing, communication, or computation nodes become unresponsive or fail. For example, coded computing techniques should be extended to the ISCC setting, accounting for e.g. sensor failures. Similarly, network coding, known for its resilience to link failures, eavesdropping, and jamming, offers a promising foundation for designing ISCC architectures capable of withstanding diverse and unpredictable network-level disruptions.
@@
Synergies of multi-tier multi-modal ISCC networks: Distributed multi-tier, multi-modal ISCC networks promise greater robustness and resilience levels through complementary features for fault tolerance. However, research in both multi-modal and distributed ISCC networks is still in its early stages, requiring new architectures and design methodologies to integrate heterogeneous devices across different locations. Moreover, fusion techniques must support efficient representation alignment such that the feature spaces across different sensor types are aligned for effective cooperative perception. Furthermore, from a multi-task perspective, multiple perception and AI tasks should be designed to share a unified feature space. This would allow a single, jointly sensed data representation to support multiple tasks, significantly reducing communication overhead.
@@
Resilience bottleneck in RBD ISCC networks: The RBD principle emphasizes embedding resilience capabilities during the design phase. In ISCC networks, resilience provisions across sensing, communication, and computing components are inherently heterogeneous and contribute differently to the end-to-end service resilience. As a result, efficient provisioning requires identifying resilience bottlenecks and strategically investing in the components that yield the greatest improvement in overall resilience. This challenge is further amplified in task-oriented designs, where optimal provisioning trade-offs depend heavily on specific task objectives. While the impact of resilience provisioning in individual subsystems has been studied, a holistic, joint approach to RBD in ISCC networks remains largely unexplored.
@@
Self-organizing ISCC networks: A critical aspect of the RBD principle is reconfigurability based on available resilience measures. However, to hugely benefit from the network reconfigurability, corresponding strategies that support effective task redistribution through resource reallocation and network reconfiguration are required. Furthermore, while existing self-organizing networks (SONs) in communication systems have partially addressed this need, they do not account for the specific requirement of ISCC networks, which involve far more complex and tightly coupled sensing, computing, and communication operations [14].
@@ V Conclusions
ISCC networks are essential for enabling intelligent perception in domains such as networked robotics, smart healthcare, Industry 5.0, and autonomous transportation. However, the interdependencies among sensing, computing, and communication components make them inherently complex and more vulnerable to uncertainties and disruptions. To address this, robustness and resilience must be embedded into their design. This article proposes an R2-ISCC framework, and outlining key principles and strategies across resource management, coding, architecture, DTs, and network-level design. A simulated case study highlights the importance and interplay of robustness and resilience. Key research problems are identified to guide future work. We hope this study inspires continued exploration and innovation toward R2-ISCC networks.

### v2

@@ IV-B Open Problems
Efficient and scalable resource management for ISCC: Although widely used in conventional networks, diversity and redundancy schemes are typically designed for individual purposes, without jointly addressing sensing, computing, and communication. In ISCC networks, new strategies that can efficiently integrate diversity and redundancy across all three components are needed, particularly for tasks that exhibit intrinsic dependencies among these functions. Moreover, the joint optimization of sensing, computing, and communication resources results in inherently high-dimensional and complex optimization problems. Hence, the development of efficient, scalable solution methods is critical to enable practical deployment and operation of robust and resilient ISCC systems.
@@
Task-oriented and criticality-aware ISCC: ISCC networks must support heterogeneous tasks with varying requirements and urgency. To improve robustness and resilience, proactive resource management must account for these task-level differences, especially during disruptions when resources are scarce and prioritization is vital. Network slicing offers a promising solution, but conventional methods focus mainly on communication. In ISCC, holistic slicing approaches that jointly manage sensing, communication, and computation resources are needed.
@@
Joint source–channel–computing network coding schemes: To support network-level resilience, coding must extend beyond individual node types and retain the network functionality even when some sensing, communication, or computation nodes become unresponsive or fail. For example, coded computing techniques should be adapted to the ISCC setting to account for failures such as missing sensor inputs. Similarly, network coding, known for its resilience to link failures, eavesdropping, and jamming, offers a promising foundation for designing ISCC architectures capable of withstanding diverse and unpredictable network-level disruptions.
@@
Synergies of multi-tier multi-modal ISCC networks: Distributed multi-tier, multi-modal ISCC networks enhance robustness and resilience through complementary fault-tolerance features. However, research in this area is still at an early stage, and require new architectures and design methodologies to integrate heterogeneous devices across locations. Effective fusion techniques are also needed to align feature spaces across sensing modalities for cooperative perception. From a multi-task perspective, designing AI and perception tasks to share a unified feature space would allow a single joint data representation to serve multiple tasks, which could greatly reduce communication overhead.
@@
Resilience bottleneck in R2BD-ISCC networks: The R2BD principle emphasizes embedding resilience capabilities during the design phase. In ISCC networks, resilience provisions across sensing, communication, and computing components are inherently heterogeneous and contribute differently to the end-to-end service resilience. Efficient provisioning requires identifying resilience bottlenecks and measuring the end-to-end resilience based on the performance metrics discussed in Section II-B, so that investments can be directed to the components that yield the greatest improvement in overall resilience. While the impact of resilience provisioning in individual subsystems has been studied, a holistic, joint approach to R2BD-ISCC networks remains largely unexplored.
@@
Self-organizing ISCC networks: A critical aspect of the R2BD principle is reconfigurability based on available resilience measures. However, to hugely benefit from the network reconfigurability, corresponding strategies that support effective task redistribution through resource reallocation and network reconfiguration are required. Furthermore, while existing self-organizing networks (SONs) in communication systems have partially addressed this need, they do not account for the specific requirement of ISCC networks, which involve far more complex and tightly coupled sensing, computing, and communication operations [15].
@@ V Conclusions
ISCC networks are essential for enabling intelligent perception in domains such as networked robotics, smart healthcare, Industry 5.0, and autonomous transportation. However, the interdependencies among sensing, computing, and communication components make them inherently complex and more vulnerable to uncertainties and disruptions. To address this, robustness and resilience must be embedded into their design. This article proposes an R2-ISCC framework, and outlining key principles and strategies across resource management, coding, DTs, architecture, and network-level design. A simulated case study highlights the importance and interplay of robustness and resilience. Key research problems are identified to guide future work. We hope this study inspires continued exploration and innovation toward R2-ISCC networks.